OpenStudy (anonymous):

if 0 is less then theta less than 360 degrees solve sec^2+2sec(2theta)=0

OpenStudy (amoodarya):



OpenStudy (amoodarya):

\[\frac{1+2\cos \theta}{\cos^2\theta}=0\\1+2\cos \theta=0\\cos \theta =\frac{-1}{2}\]

OpenStudy (anonymous):

@amoodarya can you explain where your getting the one? is that a formula?

OpenStudy (amoodarya):

\[\sec \theta =\frac{1}{\cos \theta}\]

OpenStudy (anonymous):

@amoodarya alright so how would you go about solving that then with everything applied?

OpenStudy (anonymous):

@amoodarya where would the 2sec(2theta) go in the formula? isn't it suppose to be at the bottom not top?

OpenStudy (amoodarya):

\[\sec^2\theta +2\sec \theta=0\\ \frac{ 1 }{ \cos^2\theta }+2\frac{ 1 }{ \cos \theta }=0\\\frac{ 1 }{ \cos^2\theta }+2\frac{ \cos \theta }{ \cos^2 \theta }=0\\\frac{1+2 \cos \theta }{ \cos^2 \theta }=0\\1+2 \cos \theta=0\\ \cos \theta=\frac{-1}{2}\]

OpenStudy (anonymous):

@amoodarya so the answer is -1/2 then? is there a way to check it?

OpenStudy (amoodarya):

check for what ?

OpenStudy (anonymous):

for -1/2

OpenStudy (amoodarya):

\[\cos (\theta)=-0.5\\ \theta =\pm(\frac{2\pi}{3})+2k \pi\]

OpenStudy (anonymous):

What does that mean exactly?

OpenStudy (amoodarya):

if you mean denominator goes to zero , you are in a right way but in your case , no problem \[denominator \neq 0\]

OpenStudy (amoodarya):